OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)? A. -3:2 B.-2:3 C.3:2 D.2:3 I dont understand this? Can you please help me?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

from -2 to 2 is 4 units right. this is your denominator, 4 from 4 to -2 is 6 units down. your numerator is -6 (minus, because it is down) ratio is \(-6:4\) or \(-3:2\) by removing the common factor of 2

OpenStudy (anonymous):

lol wuttt. @Hero cuz at first you say 2/3 then you say 3/2! im confused!

hero (hero):

That's because I keep making sign errors, sorry.

hero (hero):

Now, I'm agreeing with Satellite: 4 = -2m + b -2 = 2m + b b = 2m + 4 b = -2 - 2m b = b 2m + 4 = -2 - 2m 2m + 2m = -2 - 4 4m = -6 m = -6/4 m =- 3/2